Find the cube root of each of the following numbers by prime factorisation method.

 

(x) 91125

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student

(x) 91125

Solution:

91125 = 3×3×3×3×3×3×3×5×5×5

By grouping the factors in triplets of equal factors, 91125 = (3×3×3)×(3×3×3)×(5×5×5)

Here, 91125 can be grouped into triplets of equal factors,

∴ 91125 = (3×3×5) = 45

Hence, 45 is cube root of 91125.​



Thanks